OpenStudy (anonymous):

(1/64)^2/3 ( write without exponents)

OpenStudy (whpalmer4):

\[(\frac{1}{64})^{2/3}\]I would first rewrite this as \[(64^{-1})^{2/3}\]using the principle that \[a^{-n} = \frac{1}{a^n}\]Now we can apply this principle:\[(a^n)^m = a^{n*m}\]to rewrite it as \[64^{-2/3} = 64^{-1/3}*64^{-1/3}\](because exponents with a common base add when multiplying)Now we just need to find a number x such that x*x*x = 64 and we can then rewrite that as \[x^{-1}*x^{-1} = x^{-2} = \frac{1}{x^2} = \frac{1}{<whatever ~x*x =>}\]